OpenStudy (anonymous):

Eliminate the parameter. x = t + 4, y = t^2

OpenStudy (phi):

can you solve for t, using the first equation?

OpenStudy (anonymous):

I honestly have no idea what I'm doing.

OpenStudy (anonymous):

Okay so these are the answer choices: y = square root of x + x + 4 y = square root of quantity x minus four y = x2 + 16 y = x2 - 8x + 16 I'm thinking it's y = sqrt x-4 because with the first equation, t = x-4. And sqrt x-4 makes sense to me because of this. Am I completely wrong or?

OpenStudy (johnweldon1993):

Well all we do is solve for 't' in the first equation \[\large x = t + 4\] will become \[\large t = x - 4\] Now all we do is plug that into our second equation wherever we see a 't' \[\large y = t^2\] will then become \[\large y = (x - 4)^2\] And just distribute that out and you will have your answer

OpenStudy (anonymous):

So y = x^2 + 16?

OpenStudy (johnweldon1993):

Not quite You dont just apply the exponent to both terms...you expand \[\large (x-4)^2 = (x - 4)(x - 4)\]

OpenStudy (johnweldon1993):

Then you distribute \[\large x^2 - 4x - 4x + 16\] \[\large x^2 - 8x + 16\]